OpenStudy (anonymous):

find f '(0) for f(x)=(x+1)^3(4-x)^2

OpenStudy (bahrom7893):

f'(x) = (x+1)^3*2(4-x)(-1) + (4-x)^2 * 3(x+1)^2

OpenStudy (bahrom7893):

f'(0) = (0+1)^3 * 2 (4-0)(-1) + (4-0)^2 * 3(0+1)^2

OpenStudy (bahrom7893):

f'(0) = -8 + 16*3

OpenStudy (bahrom7893):

f'(0) = 48-8 = 40 (fan if I helped, thanks!)

OpenStudy (anonymous):

that helped a ton thanks